# Efficient and Scalable Graph Generation through Iterative Local Expansion

### 14 Dec 2023 v1

@@ 1 Introduction
Graphs are mathematical structures representing relational data. They comprise a set of nodes and a set of edges, denoting pairwise relations between them. This abstraction is ubiquitous in modeling discrete data across domains like social networking Fan et al. (2019), program synthesis Nguyen et al. (2012); Bieber et al. (2020), and chemical engineering Jin et al. (2018). A crucial task is the generation of new graphs possessing characteristics similar to those observed. In drug discovery, for instance, this involves creating graphs that encode the structure of a desired type of protein Ingraham et al. (2022); Martinkus et al. (2023) or molecule Jin et al. (2018); Vignac et al. (2023b).
@@
Traditional graph generation methods (Albert & Barabási, 2002) estimate parameters of known models like Stochastic Block Model (SBM)  (Holland et al., 1983) or Erdos-Renyi Erdos et al. (1960), but these often fail to capture the complexity of real-world data. Deep learning has thus become a promising alternative for graph generation, with approaches falling into two categories depending on the factorization of the data-generating distribution. Autoregressive methods build the graph sequentially by adding one node at a time and predicting the edges to the existing nodes (You et al., 2018; Liao et al., 2020; Dai et al., 2020). The need to fix the order in which the nodes are generated makes these methods sensitive to the choice of ordering and prone to overfitting. One-shot methods generate the entire graph at once using techniques like variational autoencoders (Simonovsky & Komodakis, 2018), generative adversarial networks (Cao & Kipf, 2022; Martinkus et al., 2022), normalizing flows (Liu et al., 2019), score-based, and denoising diffusion models (Niu et al., 2020; Jo et al., 2022; Haefeli et al., 2023; Vignac et al., 2023a; Yan et al., 2023).
@@
While both autoregressive and one-shot methods can generate graphs with a few hundred nodes, scaling beyond this range poses challenges. The relational structure is often modeled as a joint distribution of binary features representing edges, leading to at least quadratic complexity growth with the number of nodes. Autoregressive methods face an exponential increase in node orderings, impacting performance and potentially causing overfitting (e.g. see results of Martinkus et al. (2022)). Meanwhile, one-shot methods can struggle to capture both the global structure of the graph and the local details simultaneously, hindering their ability to generate large combinatorial structures. Finally, the neural architectures employed either exhibit limited expressiveness, although with linear complexity in the number of edges for message-passing neural networks (Xu et al., 2019; Veličković et al., 2018), or are computationally expensive with quadratic or even higher scaling factors for more expressive architectures (Dwivedi & Bresson, 2021; Maron et al., 2020).
@@
We present a novel approach to graph generation through iterative local expansion. In each step, we expand some nodes into small subgraphs and use a diffusion model to recover the appropriate local structure. For model training, we construct coarsened versions of the graphs in the dataset (Loukas, 2018; Loukas & Vandergheynst, 2018; Hermsdorff & Gunderson, 2019; Jin et al., 2020b; Kumar et al., 2022; 2023) and train a diffusion model to invert the coarsening operation as shown in Figure 1. We argue that this is inherently suitable for generating graphs, as it allows for the generation of an approximate global structure initially, followed by the addition of local details. This generative process effectively represents a particular kind of network growth, which we find to be much more robust to changes in generated graph sizes than existing approaches. Moreover, our method enables modeling the distribution of edges without the need to represent the entire joint distribution over all node pairs, enhancing scalability for larger graphs. Our theoretical analysis shows that, under mild conditions, our method exhibits subquadratic sampling complexity relative to the number of nodes for sparse graphs. We also introduce a more efficient local version of the Provably Powerful Graph Network (PPGN) (Maron et al., 2020), termed Local PPGN. This variant is especially well suited for our iterative local expansion approach, maintaining the high expressive power in the local subgraphs that we process while providing better computational efficiency. To demonstrate the effectiveness of our approach, we conducted experiments with widely used benchmark datasets. First, in the standard graph distribution modeling task from Martinkus et al. (2022); Vignac et al. (2023a), our model achieves state-of-the-art performance with the highest Validity-Uniqueness-Novelty score on the planar and tree datasets. Additionally, it generated graphs most closely matching the test set’s structural statistics for protein and point cloud datasets. Second, we evaluate our method’s ability to generalize beyond the training distribution, by generating graphs with an unseen number of nodes and verifying if they retain the defining characteristics of the training data. In this setting, our method is the only one capable of preserving these characteristics across the considered datasets. Third, we show that for sparse graphs our model exhibits subquadratic sampling complexity relative to the number of nodes, and validate this empirically by generating planar graphs of increasing size. Our implementation is available [^1] https://github.com/AndreasBergmeister/graph-generation.
@@
Figure 1: Example of a 4-level coarsening sequence. Colors indicate the node contraction sets ${\mathcal{V}}^{(p)}$. Our generation process aims at reversing with expansions and refinements the $T$ steps of this sequence from $G_{T}$ to $G_{0}$. The details of a single step are provided in Figure 2
@@ 2 Related Work
The seminal work by You et al. (2018) pioneered graph generation using recurrent neural networks, creating the adjacency matrix sequentially. Liao et al. (2020) improved this approach by simultaneously sampling edges for newly added nodes from a mixture of independent Bernoulli distributions with parameters obtained from a message-passing graph neural network. Kong et al. (2023) conceptualized this method as the inverse of an absorbing state diffusion process (Austin et al., 2021) and proposed reinforcement learning to optimize node addition sequences.
@@
Lately, diffusion models have come to dominate alternative approaches in terms of sample quality and diversity. Although initially only effective for graphs with tens of nodes (Niu et al., 2020), subsequent improvements using discrete diffusion (Vignac et al., 2023a; b; Haefeli et al., 2023), refining the diffusion process with a destination-predicting diffusion mixture (Jo et al., 2023), or dropping permutation equivariance (Yan et al., 2023) allowed for successful generation of graphs with a few hundred nodes. Nevertheless, scalability and computational complexity remain challenges for these models. As a countermeasure, Diamant et al. (2023) suggest limiting the maximal bandwidth of generated graphs. They leverage the observation that real-world graph nodes can often be ordered to confine non-zero adjacency matrix entries within a narrow diagonal band (Cuthill & McKee, 1969). Within this band, generation can be achieved using models such as GraphRNN (You et al., 2018), variational autoencoders (Grover et al., 2019), or score-based generative models (Niu et al., 2020). Alternatively, Chen et al. (2023b) introduce degree-guided diffusion, which begins with an RNN-generated degree sequence to condition the graph diffusion model. During each step, the model only considers edge connections between nodes predicted to require degree increases. This non-local process requires a simple, non-expressive, message-passing graph neural network for efficient execution. However, it does offer an increase in empirical computational efficiency.
@@
An orthogonal line of research leverages hierarchical constructions for more efficient graph generation. Dai et al. (2020) improve the original RNN-based adjacency generation by You et al. (2018) using binary tree-structured conditioning on rows and columns of the matrix, cutting complexity from $\mathcal{O}(n^{2})$ to $\mathcal{O}((n+m)\log n)$, with $n$ representing nodes and $m$ edges. Shirzad et al. (2022) suggest a two-stage process starting with tree-based cluster representation, followed by incremental subgraph construction. Another two-level approach to generation is proposed by Davies et al. (2023) using DiGress (Vignac et al., 2023a) to create cluster graphs, followed by independent generation of cluster subgraphs and intra-cluster edges. In a related vein, Karami (2023) present a methodology that extends to multiple levels of hierarchy, with autoregressive generation of cluster subgraphs. Although the independence assumptions of these hierarchical methods improve scalability, they may compromise sample accuracy, contrasting with our approach that avoids such assumptions. Both Davies et al. (2023) and Karami (2023) utilize the Louvain algorithm (Blondel et al., 2008) for pre-generating clusterings for training, unlike our method, which employs random sampling of coarsening sequences during training. Additionally, Guo et al. (2023) introduce a graph expansion layer for inclusion in the generator of a generative adversarial network or the decoder of a variational autoencoder, with parameters trained through reinforcement learning.
Hierarchical approaches have also been developed for molecular generation (Jin et al., 2019; 2020a; Kuznetsov & Polykovskiy, 2021), aiming to improve efficiency and performance by integrating domain knowledge. However, these methods are not optimized for general graph generation tasks.
@@ 3.2 Learning to Invert Graph Coarsening
An important property of this coarsening operation is that it can be inverted through an appropriate expansion and subsequent refinement step, as elaborated in Appendix A. Based on this premise, it can be deduced through an inductive argument that for any given coarsening sequence $(G=G_{0},G_{1},\dots,G_{L}=(\{v\},\emptyset))$ that transforms a graph $G$ into a single node, there exists a corresponding expansion sequence $\varpi\in\varPi(G)$ with the same elements in reverse order, i.e. $\varpi=(G_{L},\dots,G_{0})$. Note that successive coarsening steps always result in a single node graph, as long as the original graph is connected, and every coarsening step contains at least one non-trivial contraction set, i.e. a set of nodes with more than one node.
@@ 3.4 Local PPGN
A key component of our proposed methodology is the specialized architecture designed to parameterize the conditional distributions $p_{{\bm{\theta}}}({\bm{v}}_{l},{\bm{e}}_{l}\mid\tilde{G}_{l})$, or equivalently, the denoising model. Our design incorporates a novel edge-wise message-passing layer, termed Local PPGN. When designing this layer, we drew inspiration from the PPGN model (Maron et al., 2020), which is provably more expressive than graph message-passing networks at the expense of increased computational complexity (cubic in the number of nodes). Recognizing that our suggested methodology only locally alternates graphs at every expansion step and that these graphs possess a locally dense structure, as a result of the expansion process (Definition 1), we designed a layer that is locally expressive, resembles the PPGN layer on a dense (sub)graph, but retains efficiency on sparse graphs, with linear runtime relative to the number of edges. An elaborate explanation of this layer and its placement within existing graph neural network models can be found in Appendix F. In-depth architectural details of the overall model are presented in Appendix F.2.
@@ 3.7 Deterministic Expansion Size
Our graph expansion method iteratively samples a cluster size vector ${\bm{v}}$ to incrementally enlarge the graph. The process halts when ${\bm{v}}$ is entirely composed of ones, indicating no further node expansion is necessary. However, this stochastic approach may not reliably produce graphs of a predetermined size. To remedy this, we propose a deterministic expansion strategy, primarily applicable in cases of edge contraction where the maximum expansion size is two. In this strategy, ${\bm{v}}$ is treated as binary. We set the target size for the expanded graph at each expansion step and, instead of sampling ${\bm{v}}$, we we select the required number of nodes with the highest probabilities for expansion to reach the predefined size. Additionally, we introduce the reduction fraction—calculated as one minus the ratio of node counts between the original and expanded graphs—as an additional input to the model during training and inference. More details are discussed in Appendix G.
@@ 4 Experiments
Training set 0.0002 0.0310 0.0005 0.0052 0.0012 1.0 100 100
GraphRNN (You et al., 2018) 0.0049 0.2779 1.2543 0.0459 0.1034 527.4 0.0 100 100 0.0
GRAN (Liao et al., 2020) 0.0007 0.0426 0.0009 0.0075 0.0019 1.9 97.5 85.0 2.5 0.0
SPECTRE (Martinkus et al., 2022) 0.0005 0.0785 0.0012 0.0112 0.0059 2.9 25.0 100 100 25.0
DiGress (Vignac et al., 2023a) 0.0007 0.0780 0.0079 0.0098 0.0031 5.2 77.5 100 100 77.5
EDGE (Chen et al., 2023b) 0.0998 0.2672 1.0160 0.0998 0.3888 576.5 0.0 100 100 0.0
BwR (EDP-GNN) (Diamant et al., 2023) 0.0231 0.2596 0.5473 0.0444 0.1314 267.3 0.0 100 100 0.0
BiGG (Dai et al., 2020) 0.0007 0.0570 0.0367 0.0105 0.0052 17.0 62.5 85.0 42.5 5.0
Ours (one-shot) 0.0003 0.0245 0.0006 0.0104 0.0030 1.6 67.5 100 100 67.5
@@
Training set 0.0008 0.0332 0.0255 0.0063 0.0007 1.0 100 100
GraphRNN (You et al., 2018) 0.0055 0.0584 0.0785 0.0065 0.0431 14.9 5.0 100 100 5.0
GRAN (Liao et al., 2020) 0.0113 0.0553 0.0540 0.0054 0.0212 9.8 25.0 100 100 25.0
SPECTRE (Martinkus et al., 2022) 0.0015 0.0521 0.0412 0.0056 0.0028 2.0 52.5 100 100 52.5
DiGress (Vignac et al., 2023a) 0.0018 0.0485 0.0415 0.0045 0.0014 1.6 60.0 100 100 60.0
EDGE (Chen et al., 2023b) 0.2186 0.0592 0.0946 0.0952 0.1758 107.7 0.0 100 100 0.0
@@
BiGG (Dai et al., 2020) 0.0012 0.0604 0.0667 0.0059 0.0370 11.7 10.0 100 100 10.0
Ours (one-shot) 0.0141 0.0528 0.0809 0.0071 0.0205 10.4 75.0 100 100 75.0
Ours 0.0119 0.0517 0.0669 0.0067 0.0219 10.1 45.0 100 100 45.0
@@
Training set 0.0001 0.0000 0.0000 0.0127 0.0030 1.0 100 100
GRAN (Liao et al., 2020) 0.1884 0.0080 0.0199 0.2751 0.3274 601.9 0.0 100 100 0.0
DiGress (Vignac et al., 2023a) 0.0002 0.0000 0.0000 0.0113 0.0043 1.4 90.0 100 100 90.0
EDGE (Chen et al., 2023b) 0.7644 0.0000 0.0685 0.5845 0.6945 2,356.9 0.0 7.5 100 0.0
BwR (EDP-GNN) (Diamant et al., 2023) 0.0016 0.1239 0.0003 0.0480 0.0388 10.5 0.0 100 100 0.0
BiGG (Dai et al., 2020) 0.0014 0.0000 0.0000 0.0119 0.0058 5.0 100 87.5 50.0 75.0
Ours (one-shot) 0.0004 0.0000 0.0000 0.0080 0.0055 2.0 82.5 100 100 82.5
Ours 0.0001 0.0000 0.0000 0.0117 0.0047 3.8 100 100 100 100
@@
Training set 0.0003 0.0068 0.0032 0.0009 0.0003 1.0 0.0000 0.1768 0.0049 0.0064 0.0024 1.0
GraphRNN (You et al., 2018) 0.004 0.1475 0.5851 0.0152 0.0530 82.3 OOM OOM OOM OOM OOM OOM
GRAN (Liao et al., 2020) 0.0479 0.1234 0.3458 0.0125 0.0341 82.7 0.0201 0.4330 0.2625 0.0051 0.0436 18.7
SPECTRE (Martinkus et al., 2022) 0.0056 0.0843 0.0267 0.0052 0.0118 16.9 OOM OOM OOM OOM OOM OOM
DiGress (Vignac et al., 2023a) 0.0041 0.0489 0.1286 0.0018 0.0065 17.0 OOM OOM OOM OOM OOM OOM
EDGE (Chen et al., 2023b) 0.5131 0.0825 0.7326 0.3044 0.6395 884.3 0.5398 0.8201 1.0730 0.6084 0.7339 155.2
BwR (EDP-GNN) (Diamant et al., 2023) 0.1262 0.4202 0.4939 0.0702 0.1199 222.9 0.4927 0.4690 1.0730 0.2912 0.5916 127.8
BiGG (Dai et al., 2020) 0.0070 0.1150 0.4696 0.0067 0.0222 53.7 0.0994 0.6035 0.3633 0.1589 0.0994 35.8
Ours (one-shot) 0.0015 0.0711 0.0396 0.0026 0.0086 11.9 OOM OOM OOM OOM OOM OOM
Ours 0.0030 0.0309 0.0047 0.0013 0.0030 5.5 0.0139 0.5775 0.0780 0.0055 0.0186 6.9
Table 2: Sample quality on real-world graphs. All models achieved perfect uniqueness and novelty.
@@
In Table 1 the most critical metric is the percentage of valid, unique, and novel graphs (V.U.N) in the generated set. Validity for synthetic graphs tests if the core property of the graph is respected, e.g. if the generated graph is planar or if the generated tree has no cycles. Uniqueness and novelty metrics report the diversity of the output, serving as an indicator for non-overfitting. It can be seen that our approach performs well and outperforms our baseline, which does not use coarsening but directly runs our diffusion model on the full graph (Ours (one-shot)). The only dataset where this is not the case is the SBM dataset, and this is likely due to the fact that SBM graphs have limited structure. Aside from the large clusters, the graphs are essentially random. Note that our model still achieves a satisfactory V.U.N. score.
The first five columns show the maximum mean discrepancy (MMD) between the generated and training graphs for the degree distribution, clustering coefficient, orbit counts, spectrum, and wavelet coefficients. These metrics are summarized as the mean ratio of these statistical similarities compared to those achieved by the training set. Although DiGress is the overall best performer with respect to this metric, our method achieves competitive results and is the best for planar graphs.
The benefits of our approach become clearer with larger and more complicated real-world graphs. In Table 2 we show model performance on protein graphs with up to 500 nodes and point cloud graphs with up to 5037 nodes (see Table 6). In these cases, our model improves by a factor of more than 3 the MMD ratio over the competitors. Note that several methods are unable to scale to up to 5037 nodes.
A comprehensive discussion on the sample complexity of our method, along with a comparison of empirical runtimes against other methods, can be found in Appendix H. Sample graphs generated by our model are presented in Appendix J.

### 21 Feb 2024 v2

@@ 1 Introduction
Graphs are mathematical structures representing relational data. They comprise a set of nodes and a set of edges, denoting pairwise relations between them. This abstraction is ubiquitous in modeling discrete data across domains like social networking (Fan et al., 2019), program synthesis (Nguyen et al., 2012; Bieber et al., 2020), or even origami design (Geiger et al., 2023). A crucial task is the generation of new graphs that possess characteristics similar to those observed. For example, in drug discovery, this involves creating graphs that encode the structure of a desired type of protein (Ingraham et al., 2022; Martinkus et al., 2023) or molecule (Jin et al., 2018; Vignac et al., 2023b).
@@
Traditional graph generation methods (Albert & Barabási, 2002) estimate parameters of known models like Stochastic Block Model (SBM)  (Holland et al., 1983) or Erdos-Renyi Erdos et al. (1960), but often fail to capture the complexity of real-world data. Deep learning offers a promising alternative, with approaches falling into two categories depending on the factorization of the data-generating distribution. Autoregressive techniques build graphs incrementally, predicting edges for each new node (You et al., 2018; Liao et al., 2020; Dai et al., 2020). One-shot methods generate the entire graph at once using techniques such as variational autoencoders (Simonovsky & Komodakis, 2018), generative adversarial networks (Cao & Kipf, 2022; Martinkus et al., 2022), normalizing flows (Liu et al., 2019), score-based and denoising diffusion models (Niu et al., 2020; Haefeli et al., 2023).
@@
Despite the success of these methods in generating graphs comprising several hundred nodes, scaling beyond this range poses challenges. The computational cost of predicting edges between all node pairs scales at least quadratically with the number of nodes, which is inefficient for sparse graphs typical of real-world data. Sample fidelity is also an issue, as autoregressive methods struggle with node-permutation-invariant training due to the factorial increase in node orderings, and one-shot methods often fail to capture both global and local graph structure simultaneously. Also, in contrast to algorithmic approaches (Babiac et al., 2023), neither have been shown to generalize to larger unseen graph sizes. Finally, the neural architectures employed either exhibit limited expressiveness, although with linear complexity in the number of edges for message passing neural networks (Xu et al., 2019), or are computationally expensive with quadratic or even higher scaling factors for more expressive architectures (Dwivedi & Bresson, 2021; Maron et al., 2020).
@@
We present a novel approach to graph generation through iterative local expansion. In each step, we expand some nodes into small subgraphs and use a diffusion model to recover the appropriate local structure. The model is trained to reverse a graph coarsening process, as depicted in Figure 1, applied to the dataset graphs (Loukas, 2018; Loukas & Vandergheynst, 2018; Hermsdorff & Gunderson, 2019; Jin et al., 2020b; Kumar et al., 2022; 2023). We argue that this is inherently suitable for generating graphs, as it allows for the generation of an approximate global structure initially, followed by the addition of local details. This generative process effectively represents a particular kind of network growth, which we find to be much more robust to changes in generated graph sizes than existing approaches. Moreover, our method enables modeling the distribution of edges without the need to represent the entire joint distribution over all node pairs, enhancing scalability for larger graphs. Our theoretical analysis shows that, under mild conditions, our method exhibits subquadratic sampling complexity relative to the number of nodes for sparse graphs. We also introduce a more efficient local version of the Provably Powerful Graph Network (PPGN) (Maron et al., 2020), termed Local PPGN. This variant is especially well suited for our iterative local expansion approach, maintaining the high expressive power in the local subgraphs that we process while providing better computational efficiency. To demonstrate the effectiveness of our approach, we conducted experiments with widely used benchmark datasets. First, in the standard graph distribution modeling task from Martinkus et al. (2022); Vignac et al. (2023a), our model achieves state-of-the-art performance with the highest Validity-Uniqueness-Novelty score on the planar and tree datasets. Additionally, it generated graphs most closely matching the test set’s structural statistics for protein and point cloud datasets. Second, we evaluate our method’s ability to generalize beyond the training distribution, by generating graphs with an unseen number of nodes and verifying if they retain the defining characteristics of the training data. In this setting, our method is the only one capable of preserving these characteristics across the considered datasets. Third, we show that for sparse graphs our model exhibits subquadratic sampling complexity relative to the number of nodes, and validate this empirically by generating planar graphs of increasing size. Our implementation is available [^1] https://github.com/AndreasBergmeister/graph-generation.
@@
Figure 1: Example of a 4-level coarsening sequence. Colors indicate the node contraction sets ${\mathcal{V}}^{(p)}$. Our generation process aims at reversing with expansions and refinements the $T$ steps of this sequence from $G_{T}$ to $G_{0}$. The details of a single step are provided in Figure 2.
@@ 2 Related Work
The seminal work by You et al. (2018) pioneered graph generation using recurrent neural networks, creating the adjacency matrix sequentially. Liao et al. (2020) improved this approach by simultaneously sampling edges for newly added nodes from a mixture of independent Bernoulli distributions with parameters obtained from a message passing graph neural network. Kong et al. (2023) conceptualized this method as the inverse of an absorbing state diffusion process (Austin et al., 2021) and proposed reinforcement learning to optimize node addition sequences.
@@
Lately, diffusion models have come to dominate alternative approaches in terms of sample quality and diversity. Although initially only effective for graphs with tens of nodes (Niu et al., 2020), subsequent improvements using discrete diffusion (Vignac et al., 2023a; b; Haefeli et al., 2023), refining the diffusion process with a destination-predicting diffusion mixture (Jo et al., 2023), or dropping permutation equivariance (Yan et al., 2023) allowed for successful generation of graphs with a few hundred nodes. Nevertheless, scalability and computational complexity remain challenges for these models. As a countermeasure, Diamant et al. (2023) suggest limiting the maximal bandwidth of generated graphs. They leverage the observation that real-world graph nodes can often be ordered to confine non-zero adjacency matrix entries within a narrow diagonal band (Cuthill & McKee, 1969). Within this band, generation can be achieved using models such as GraphRNN (You et al., 2018), variational autoencoders (Grover et al., 2019), or score-based generative models (Niu et al., 2020). Alternatively, Chen et al. (2023b) introduce degree-guided diffusion, which begins with an RNN-generated degree sequence to condition the graph diffusion model. During each step, the model only considers edge connections between nodes predicted to require degree increases. This non-local process requires a simple, non-expressive, message passing graph neural network for efficient execution. However, it does offer an increase in empirical computational efficiency. Goyal et al. (2020) propose a different approach by generating a canonical string representation of the graph using a long short-term memory network. Although the length of the string is linear in the number of graph edges, generating the strings for model training has worst-case factorial complexity, which limits the practicality of this approach for general large-scale graph generation tasks.
@@
An orthogonal line of research leverages hierarchical constructions for more efficient graph generation. Dai et al. (2020) improve the original RNN-based adjacency generation by You et al. (2018) using binary tree-structured conditioning on rows and columns of the matrix, cutting the complexity from $\mathcal{O}(n^{2})$ to $\mathcal{O}((n+m)\log n)$, with $n$ representing nodes and $m$ edges. Shirzad et al. (2022) suggest a two-stage process starting with tree-based cluster representation, followed by incremental subgraph construction. Another two-level approach to generation is proposed by Davies et al. (2023) using DiGress (Vignac et al., 2023a) to create cluster graphs, followed by independent generation of cluster subgraphs and intra-cluster edges. In a related vein, Karami (2023) present a methodology that extends to multiple levels of hierarchy, with autoregressive generation of cluster subgraphs. Limnios et al. (2023) propose another method to enhance DiGress’s scalability, which involves a divide-and-conquer strategy for sampling subgraph coverings. Although the independence assumptions of these hierarchical methods improve scalability, they may compromise sample accuracy, contrasting with our approach that avoids such assumptions. Both Davies et al. (2023) and Karami (2023) utilize the Louvain algorithm (Blondel et al., 2008) for pre-generating clusterings for training, unlike our method, which employs random sampling of coarsening sequences during training. Additionally, Guo et al. (2023) introduce a graph expansion layer for inclusion in the generator of a generative adversarial network or the decoder of a variational autoencoder, with parameter training carried out through reinforcement learning. Hierarchical approaches have also been developed for molecular generation (Jin et al., 2018; 2020a; Kuznetsov & Polykovskiy, 2021), with the aim of improving efficiency and performance by integrating domain knowledge. However, these methods are not optimized for general graph generation tasks.
@@ 3.2 Learning to Invert Graph Coarsening
An important property of this coarsening operation is that it can be inverted through an appropriate expansion and subsequent refinement step, as elaborated in Appendix A. Based on this premise, it can be deduced through an inductive argument that for any given coarsening sequence $(G=G_{0},G_{1},\dots,G_{L}=(\{v\},\emptyset))$ that transforms a graph $G$ into a single node, there exists a corresponding expansion sequence $\varpi\in\varPi(G)$ with the same elements in reverse order, i.e. $\varpi=(G_{L},\dots,G_{0})$. Note that successive coarsening steps always result in a single-node graph, as long as the original graph is connected, and every coarsening step contains at least one non-trivial contraction set, i.e. a set of nodes with more than one node.
@@ 3.4 Local PPGN
A key component of our proposed methodology is the specialized architecture designed to parameterize the conditional distributions $p_{{\bm{\theta}}}({\bm{v}}_{l},{\bm{e}}_{l}\mid\tilde{G}_{l})$, or equivalently, the denoising model. Our design incorporates a novel edge-wise message passing layer, termed Local PPGN. When designing this layer, we drew inspiration from the PPGN model (Maron et al., 2020), which is provably more expressive than graph message passing networks at the expense of increased computational complexity (cubic in the number of nodes). Recognizing that our suggested methodology only locally alternates graphs at every expansion step and that these graphs possess a locally dense structure, as a result of the expansion process (Definition 1), we designed a layer that is locally expressive, resembles the PPGN layer on a dense (sub)graph, but retains efficiency on sparse graphs, with linear runtime relative to the number of edges. An elaborate explanation of this layer and its placement within existing graph neural network models can be found in Appendix F. In-depth architectural details of the overall model are presented in Appendix F.2.
@@ 3.7 Deterministic Expansion Size
Our graph expansion method iteratively samples a cluster size vector ${\bm{v}}$ to incrementally enlarge the graph. The process halts when ${\bm{v}}$ is entirely composed of ones, indicating no further node expansion is necessary. However, this stochastic approach may not reliably produce graphs of a predetermined size. To remedy this, we propose a deterministic expansion strategy, primarily applicable in cases of edge contraction where the maximum expansion size is two. In this strategy, ${\bm{v}}$ is treated as binary. We set the target size for the expanded graph at each expansion step and, instead of sampling ${\bm{v}}$, we select the required number of nodes with the highest probabilities for expansion to reach the predefined size. Additionally, we introduce the reduction fraction, calculated as one minus the ratio of node counts between the original and expanded graphs, as an additional input to the model during training and inference. More details are discussed in Appendix G.
@@ 4 Experiments
Training set 0.0002 0.0310 0.0005 0.0038 0.0012 1.0 100 100
GraphRNN (You et al., 2018) 0.0049 0.2779 1.2543 0.0459 0.1034 490.2 0.0 100 100 0.0
GRAN (Liao et al., 2020) 0.0007 0.0426 0.0009 0.0075 0.0019 2.0 97.5 85.0 2.5 0.0
SPECTRE (Martinkus et al., 2022) 0.0005 0.0785 0.0012 0.0112 0.0059 3.0 25.0 100 100 25.0
DiGress (Vignac et al., 2023a) 0.0007 0.0780 0.0079 0.0098 0.0031 5.1 77.5 100 100 77.5
EDGE (Chen et al., 2023b) 0.0761 0.3229 0.7737 0.0957 0.3627 431.4 0.0 100 100 0.0
BwR (EDP-GNN) (Diamant et al., 2023) 0.0231 0.2596 0.5473 0.0444 0.1314 251.9 0.0 100 100 0.0
BiGG (Dai et al., 2020) 0.0007 0.0570 0.0367 0.0105 0.0052 16.0 62.5 85.0 42.5 5.0
GraphGen (Goyal et al., 2020) 0.0328 0.2106 0.4236 0.0430 0.0989 210.3 7.5 100 100 100
Ours (one-shot) 0.0003 0.0245 0.0006 0.0104 0.0030 1.7 67.5 100 100 67.5
@@
Training set 0.0008 0.0332 0.0255 0.0027 0.0007 1.0 100 100
GraphRNN (You et al., 2018) 0.0055 0.0584 0.0785 0.0065 0.0431 14.7 5.0 100 100 5.0
GRAN (Liao et al., 2020) 0.0113 0.0553 0.0540 0.0054 0.0212 9.7 25.0 100 100 25.0
SPECTRE (Martinkus et al., 2022) 0.0015 0.0521 0.0412 0.0056 0.0028 2.2 52.5 100 100 52.5
DiGress (Vignac et al., 2023a) 0.0018 0.0485 0.0415 0.0045 0.0014 1.7 60.0 100 100 60.0
EDGE (Chen et al., 2023b) 0.0279 0.1113 0.0854 0.0251 0.1500 51.4 0.0 100 100 0.0
@@
BiGG (Dai et al., 2020) 0.0012 0.0604 0.0667 0.0059 0.0370 11.9 10.0 100 100 10.0
GraphGen (Goyal et al., 2020) 0.0550 0.0623 0.1189 0.0182 0.1193 48.8 5.0 100 100 5.0
Ours (one-shot) 0.0141 0.0528 0.0809 0.0071 0.0205 10.5 75.0 100 100 75.0
Ours 0.0119 0.0517 0.0669 0.0067 0.0219 10.2 45.0 100 100 45.0
@@
Training set 0.0001 0.0000 0.0000 0.0075 0.0030 1.0 100 100
GRAN (Liao et al., 2020) 0.1884 0.0080 0.0199 0.2751 0.3274 607.0 0.0 100 100 0.0
DiGress (Vignac et al., 2023a) 0.0002 0.0000 0.0000 0.0113 0.0043 1.6 90.0 100 100 90.0
EDGE (Chen et al., 2023b) 0.2678 0.0000 0.7357 0.2247 0.4230 850.7 0.0 7.5 100 0.0
BwR (EDP-GNN) (Diamant et al., 2023) 0.0016 0.1239 0.0003 0.0480 0.0388 11.4 0.0 100 100 0.0
BiGG (Dai et al., 2020) 0.0014 0.0000 0.0000 0.0119 0.0058 5.2 100 87.5 50.0 75.0
GraphGen (Goyal et al., 2020) 0.0105 0.0000 0.0000 0.0153 0.0122 33.2 95.0 100 100 95.0
Ours (one-shot) 0.0004 0.0000 0.0000 0.0080 0.0055 2.1 82.5 100 100 82.5
Ours 0.0001 0.0000 0.0000 0.0117 0.0047 4.0 100 100 100 100
@@
Training set 0.0003 0.0068 0.0032 0.0005 0.0003 1.0 0.0000 0.1768 0.0049 0.0043 0.0024 1.0
GraphRNN (You et al., 2018) 0.004 0.1475 0.5851 0.0152 0.0530 91.3 OOM OOM OOM OOM OOM OOM
GRAN (Liao et al., 2020) 0.0479 0.1234 0.3458 0.0125 0.0341 87.5 0.0201 0.4330 0.2625 0.0051 0.0436 18.8
SPECTRE (Martinkus et al., 2022) 0.0056 0.0843 0.0267 0.0052 0.0118 19.0 OOM OOM OOM OOM OOM OOM
DiGress (Vignac et al., 2023a) 0.0041 0.0489 0.1286 0.0018 0.0065 18.0 OOM OOM OOM OOM OOM OOM
EDGE (Chen et al., 2023b) 0.1863 0.3406 0.6786 0.1075 0.2371 399.1 0.4441 0.3298 1.0730 0.4006 0.6310 143.4
BwR (EDP-GNN) (Diamant et al., 2023) 0.1262 0.4202 0.4939 0.0702 0.1199 245.4 0.4927 0.4690 1.0730 0.2912 0.5916 133.2
BiGG (Dai et al., 2020) 0.0070 0.1150 0.4696 0.0067 0.0222 57.5 0.0994 0.6035 0.3633 0.1589 0.0994 38.8
GraphGen (Goyal et al., 2020) 0.0159 0.1677 0.3789 0.0181 0.0477 83.5 OOT OOT OOT OOT OOT OOT
Ours (one-shot) 0.0015 0.0711 0.0396 0.0026 0.0086 13.3 OOM OOM OOM OOM OOM OOM
Ours 0.0030 0.0309 0.0047 0.0013 0.0030 5.9 0.0139 0.5775 0.0780 0.0055 0.0186 7.0
Table 2: Sample quality on real-world graphs. All models achieve perfect uniqueness and novelty. Several models fail on the point cloud dataset due to memory limitations (OOM), and GraphGen is unable to generate the canonical string representations within a reasonable timeframe (OOT).
@@
In Table 1 the most critical metric is the percentage of valid, unique, and novel graphs (V.U.N) in the generated set. Validity for synthetic graphs indicates the adherence to the defined properties, e.g. planarity or acyclicity. Uniqueness and novelty metrics report the diversity of the output, serving as an indicator for non-overfitting. Our method demonstrates strong performance, surpassing our baseline, which operates without iterative expansion, but directly generates the full graph using the diffusion model (Ours (one-shot)). The exception is the SBM dataset, where the inherent randomness of the graphs and the absence structure aside from large clusters, likely affects the results. Nevertheless, our model still attains a satisfactory V.U.N. score. The first five columns of the table show the maximum mean discrepancy (MMD) between the generated and test graphs for the degree distribution, clustering coefficient, orbit counts, spectrum, and wavelet coefficients. We summarize these metrics by the average ratio between the generated and training MMDs. Although DiGress is the overall best performer with respect to this metric, our method achieves competitive results and is the best for planar graphs.
The benefits of our approach become clearer with larger, complex real-world graphs. In Table 2 we show model performance on protein graphs with up to 500 nodes and point cloud graphs with up to 5037 nodes (see Table 6). In both cases, our method outperforms competitors by a large margin in structural similarity to the test set. Note that several methods are unable to scale to 5037 nodes.
Appendix H offers a runtime comparison, affirming our method’s subquadratic scaling when generating sparse graphs of increasing size. This section also includes a theoretical analysis of the model’s complexity. Sample graphs generated by our model can be found in Appendix J.